3w
____ + 8 = w - 7 ____
4 3



Answer by stanbon(57246)

3w
____ + 8 = w - 7 ____
4 3
==========================
(3/4)w + 8 = w - (7/3)
-----
Multiply thru by 12:
9w + 72 = 12w - 28
--------
3w = 100
--------
w = 33 1/3
=================
Cheers,
Stan H.
=================